OpenStudy (anonymous):

find the limit or show that it does not exist. lim (x-4)/(x^2+7) x--> infinity

OpenStudy (anonymous):

degree of the denominator is larger than the degree of the numerator so limit is 0

OpenStudy (anonymous):

In general, if the degree of the polynomial on the top is less than that on the bottom, then the limit is 0. Try that with l'Hopital's Rule. Beaten!

OpenStudy (anonymous):

this is identical to a pre calc question " find the horizontal asymptote if any" and the answer would be \(y=0\)

OpenStudy (anonymous):

think of what would happen if you replaced \(x\) by \(1000000=10^6\) numerator would have 6 places, but the denominator would have 12!

OpenStudy (anonymous):

ok. so determining the infinite limits are the same as finding the horizontal asymptote? And the vertical asymptotes are not related in this way?

OpenStudy (anonymous):

good explanations, thank you

OpenStudy (anonymous):

yes, it is the same

OpenStudy (anonymous):

well, it is the same if you are looking for a limit as x goes to infinity not for example as x goes to zero

OpenStudy (anonymous):

and if there is a square root in the expression does the same rule apply?

OpenStudy (anonymous):

so you can use the same rules you knew from whatever pre calc class you had

OpenStudy (anonymous):

essentially for example \[\lim_{x\to \infty}\frac{5x}{\sqrt{6x^2+x+4}}=\frac{5}{\sqrt{6}}\]

OpenStudy (anonymous):

ok, so the limit would still be 0 because the denominator has a larger degree than the numerator

OpenStudy (anonymous):

thank you

OpenStudy (anonymous):

in my example the limit was \(\frac{5}{\sqrt{6}}\) because \(\sqrt{6x^2}\) is like \[\sqrt{6}x\] so the degrees are the same, both 1

OpenStudy (anonymous):

ok, I was thinking you had to take out the entire expression out of the radical in order for them to be the same